A sample of 150 people from a certain industrial community showed that 80 people suffered from a lung disease. A sample of 100 people from a rural community showed that 30 suffered from the same lung disease. At α =0.05, is there a difference between the proportions of people who suffer from the disease in the two communities?
1
Expert's answer
2020-06-09T18:05:48-0400

"\text{Let the proportion of industrial community is}\; p_{1}\\
\text{and the proportion of rural community is}\; p_{2},\\
\hat p_{1}=\frac{80}{150}=0.533\\
\hat p_{2}=\frac{30}{100}=0.3\\
\hat p_{p}=\frac{80+30}{150+100}=0.44\\
H_0: p_{1}= p_{2}\\
H_1:p_{1}\neq p_{2}\\
Let \; \alpha =0.05\implies Z_{\frac{α}{2}}=Z_{0.025}=1.96,\\
Z=\frac{\hat p_{1} -\hat p_{2}}{\sqrt{\hat p_{p}(1-\hat p_{p})(\frac{1}{n_{1}}+\frac{1}{n_{2}})}}\\
Z=\frac{0.533 -0.3}{\sqrt{0.44(0.56)(\frac{1}{150}+\frac{1}{100})}}=3.64\\
\text{the rejection region} :\;z>1.96, z<-1.96\\
\text{The decision is to reject}:\; H_0\\
\text{this means that there is a difference between }\\
\text{ the proportions of people who suffer from }\\
\text{the disease in the two communities }"